# Counting Stacked Objects

### v3

@@ 3.2 Occupancy Ratio Estimation
To create it, we used the ABC dataset [8] that features a wide variety of computer-assisted design (CAD) models. We retained only watertight objects with a single connected component and rescaled them to fit in a cube of side $0.05$. We then generated a virtual 3D scene with a container, and used a physics-based simulator to drop an initial batch of 100 identical objects in that box. We repeated this step a random number of times or until the box was full, that is to say, the union of objects reached the space above the box after the physical simulation has converged. In each scene, the container is given a random shape and scale. We also include some scenes without any container where objects are directly stacked on the floor, and some scenes where boxes are partially full, as is often the case inreal-life.
@@ 4.3 Occupancy Ratio Evaluation
This first baseline tends to predict values lower than expected. Thus, we defined a second one we dubbed DepthCorrected. It uses linear regression to correct $\gamma_{est}^{norm}$ into a new estimate $\gamma_{est}^{corrected}$. This yields a method able to model the observation that depth maps with high variance tend to correspond to low occupancy ratios. Finally, we also compare with the mean estimator, that predicts the mean percentage of $32.3\%$ occupancy ratio for all inputs.
@@ 5 Conclusion
We would like to sincerely thank all the participants who spent time to play our counting game and provided us with valuable data on human performance at this task. In particular, we would like to congratulate Martin Engilberge, who gave the most accurate estimates and has been awarded a box of chocolates. We would also like to thank Nicolas Talabot, Víctor Batlle and Adriano D’Alessandro for insightful discussions.
@@ 7 Dataset Details
We drop batches of objects, arranged in a $4\times 4\times 5$ grid, into a box positioned at $(0,0,0.5)$. The box has a side length of $1$ and a thickness of $0.04$. Once the simulation converges, we check if the union of the objects intersects with an invisible cube placed directly on top of the box. If an intersection occurs, the simulation stops, and objects outside the box are deleted. If no intersection is detected, a new batch of objects is added, and the simulation is performed again.
@@
Finally, we export the calibrated camera parameters in a format compatible with nerfstudio [25]. Since these cameras are not produced by COLMAP, they do not include a 3D point cloud that 3DGS can use as initialization. This poses a challenge, as a fully random initialization may generate distant Gaussian points outside the cameras’ range, which are not removed and interfere with the volume estimation. To address this, we generate a set of 100 grey points within the unit cube, centered at $(0,0,0.5)$. This simple initialization proves sufficient to quickly produce a faithful 3D reconstruction and resolves the aforementioned issue.
@@ 12 Architecture details
To predict a scalar value from the $32\times 32\times 768$ feature image produced by DinoV2, we employ a series of convolutional layers to progressively reduce both the spatial dimensions and the number of channels. The convolutional layers successively reduce the channel dimension from the initial 768 down to 512, 256, 128, and finally 64. Concurrently, the spatial dimensions of the feature map are reduced from $32\times 32$ to $16\times 16$, $8\times 8$, $4\times 4$, and ultimately $2\times 2$.
Following this, an adaptive average pooling layer compresses the spatial dimensions to a single pixel while preserving the 64-channel depth. The resulting $1\times 1\times 64$ tensor is passed through a fully connected linear layer to map it to a scalar output. Finally, a sigmoid activation function is applied to produce the final prediction in the $[0,1]$ range.

### v4

@@ 3.2 Occupancy Ratio Estimation
To create it, we used the ABC dataset [8] that features a wide variety of computer-assisted design (CAD) models. We retained only watertight objects with a single connected component and rescaled them to fit in a cube of side 0.050.050.05. We then generated a virtual 3D scene with a container, and used a physics-based simulator to drop an initial batch of 100 identical objects in that box. We repeated this step a random number of times or until the box was full, that is to say, the union of objects reached the space above the box after the physical simulation has converged. In each scene, the container is given a random shape and scale. We also include some scenes without any container where objects are directly stacked on the floor, and some scenes where boxes are partially full, as is often the case inreal-life.
@@ 4.3 Occupancy Ratio Evaluation
[…]
@@ 5 Conclusion
We would like to sincerely thank all the participants who spent time to play our counting game and provided us with valuable data on human performance at this task. In particular, we would like to congratulate Martin Engilberge, who gave the most accurate estimates and has been awarded a box of chocolates. We would also like to thank Nicolas Talabot, Víctor Batlle and Adriano D’Alessandro for insightful discussions. This work was supported in part by the Swiss National Science Foundation.
@@ 7 Dataset Details
We drop batches of objects, arranged in a 4×4×54$\times$ 4$\times$ 54 × 4 × 5 grid, into a box positioned at (0,0,0.5)(0,0,0.5)( 0 , 0 , 0.5 ). The box has a side length of 111 and a thickness of 0.040.040.04. Once the simulation converges, we check if the union of the objects intersects with an invisible cube placed directly on top of the box. If an intersection occurs, the simulation stops, and objects outside the box are deleted. If no intersection is detected, a new batch of objects is added, and the simulation is performed again.
@@
Finally, we export the calibrated camera parameters in a format compatible with nerfstudio [25]. Since these cameras are not produced by COLMAP, they do not include a 3D point cloud that 3DGS can use as initialization. This poses a challenge, as a fully random initialization may generate distant Gaussian points outside the cameras’ range, which are not removed and interfere with the volume estimation. To address this, we generate a set of 100 grey points within the unit cube, centered at (0,0,0.5)(0,0,0.5)( 0 , 0 , 0.5 ). This simple initialization proves sufficient to quickly produce a faithful 3D reconstruction and resolves the aforementioned issue.
@@ 12 Architecture details
To predict a scalar value from the 32×32×76832$\times$ 32$\times$ 76832 × 32 × 768 feature image produced by DinoV2, we employ a series of convolutional layers to progressively reduce both the spatial dimensions and the number of channels. The convolutional layers successively reduce the channel dimension from the initial 768 down to 512, 256, 128, and finally 64. Concurrently, the spatial dimensions of the feature map are reduced from $\times 32$ 32 to $\times 16$ 16, $\times 8$ 8, $\times 4$ 4, and ultimately $\times 2$ 2.
Following this, an adaptive average pooling layer compresses the spatial dimensions to a single pixel while preserving the 64-channel depth. The resulting 1×1×641$\times$ 1$\times$ 641 × 1 × 64 tensor is passed through a fully connected linear layer to map it to a scalar output. Finally, a sigmoid activation function is applied to produce the final prediction in the [0,1][0,1][ 0 , 1 ] range.